Need help please.
5) Suppose you travel north for 65 kilometers then travel east 75 kilometers. How far are you from your starting point?
Answer:
Show your work in this space.

Answer by fractalier(6550)

If you make a quick diagram, you discover that the distance from your starting point is just the hypotenuse of a right triangle whose legs are 65 and 75 km.
Using the Pythagorean Theorem, we have
c^2 = a^2 + b^2
c^2 = 65^2 + 75^2
c^2 = 4225 + 5625
c^2 = 9850 and
c = square root of 9850 or about 99.25 km